OpenStudy (andijo76):

If the odds of winning a raffle are 18:208, what is the probability of winning? 18/208 18/226 208/18 208/226

OpenStudy (ash2326):

odds in favour= 18/208 no. of ways in which one can win raffle=18 no. of ways of not winning=208 total no. of ways=208+18=226 Probability of winning= (no. of ways of winning)/(total no. of ways)=18/226